"A baseball team played 5 more games this season than last season. Last season, the team won 60% of their games, and this season it won 55% of their games. The team won the same number of games this season as last season. Find the number of games played each season."
I AM NOT SURE HOW TO GO ABOUT SOLVING THIS...IS THERE A FORMULA TO FOLLOW? THANKS TONS!

Answer by stanbon(75887)

"A baseball team played 5 more games this season than last season. Last season, the team won 60% of their games, and this season it won 55% of their games. The team won the same number of games this season as last season. Find the number of games played each season."
I AM NOT SURE HOW TO GO ABOUT SOLVING THIS...IS THERE A FORMULA TO FOLLOW? THANKS TONS!
-----
Let # of games played last season be "x".
Team won 0.6x of the games.
-----------------------
Then # of games played this season is "x+5".
Team won 0.55(x+5) of the games.
-----
Equation:
0.6x = 0.55(x+5)
60x = 55x + 55*5
5x = 55*5
x = 55 (# of games they played last season)
x+5 = 60 (# of games they played this season)
=========================
Cheers,
Stan H.
==============